# Towards shutdownable agents via stochastic choice

### v6

@@ Appendix D Proof that DReST-optimal policies are maximally USEFUL and maximally NEUTRAL
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\\ \mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)\\ +\sum_{i=1}^{s-1}\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1-i)\\ -\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-i)\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (1)
Here is how to interpret this equation. Selecting trajectory-length $x$ in mini-episode $e_{s}$ is guaranteed to yield at least $\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)$: the expected return that would be had if $x$ were selected in all $s-1$ previous mini-episodes. In addition, there is a probability of $Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-2\}$ that selecting $x$ in $e_{s}$ yields $\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-2)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)\big{)}$: the extra expected return that would be had if $x$ were selected in only $s-2$ previous mini-episodes. In addition, there is a probability of $Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-3\}$ that selecting $x$ in $e_{s}$ yields $\big{(}\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-3)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-2)\big{)}$: the extra expected return that would be had if $x$ were selected in only $s-3$ previous mini-episodes. And so on.
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (2)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=y)=\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}$ (3)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x\lor L=y)=\\ Pr_{\pi,e_{s}}\{L=x\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}\bigg{)}\\ +Pr_{\pi,e_{s}}\{L=y\}\cdot\bigg{(}\lambda^{s-1-\frac{s-1}{k}}\\ +\sum_{i=1}^{s-1}\big{(}\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big{)}\\ \cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}\bigg{)}$ (4)
@@
(Note that $\pi_{n-1}$ implies one probability distribution over trajectory-lengths in the first $n-1$ mini-episodes $e_{1}$ to $e_{n-1}$ and implies a different probability distribution over trajectory-lengths in the final mini-episode $e_{n}$. Given that the environments in mini-episodes $e_{1}$ to $e_{n}$ are observationally-equivalent, policies like $\pi_{n-1}$ cannot be implemented. Nevertheless, it is useful to refer to policies like $\pi_{n-1}$ in proving Lemma D.2.)
Let $\pi_{n-2}$ be identical to $\pi_{n}$ except that $\pi_{n-2}$ selects trajectory-lengths $x$ and $y$ with the same probability $\mu$ in the final two mini-episodes $e_{n-1}$ to $e_{n}$. More precisely, $\pi_{n-2}$ is such that $Pr_{\pi_{n-2},e_{n-1}-e_{n}}\{L=x\}=Pr_{\pi_{n-2},e_{n-1}-e_{n}}\{L=y\}=\mu$.
And so on.
@@
$\mathbb{E}_{\pi_{a-1},E}(R|L=x\lor L=y)=\\ \mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (5)
@@
$\mathbb{E}_{\pi_{a},E}(R|L=x\lor L=y)=\\ \mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (6)
@@
$\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}+\\ \sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\\ >\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg{)}\\ +(\mu-\Delta)\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}\\ +\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg{)}\$
$+\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg{)}\\ +\mu\cdot\bigg{(}\lambda^{j-\frac{j}{k}}\\ +\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg{)}\bigg{)}\$ (7)
@@
$\Delta\cdot\bigg{(}\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big{(}\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big{)}\\ \cdot(Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\\ -Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\})\bigg{)}\\ +\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\\ +Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}>0$ (8)
By stipulation, $\Delta>0$. And since $10<\lambda<$ 1, $\lambda^{a-1-\frac{a-1}{k}}>0$ and $\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}>0$ for all $a$, $n$, and $k$. And since $Pr_{\pi_{a},e_{1}-e_{a}}\{L=x\}>Pr_{\pi_{a},e_{1}-e_{a}}\{L=y\}$, $Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}--Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\geq 0$ for all $a$ and $i$ and $Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}--Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}>0$ for all $a$ and some $i$ such that $11\leq i\leq a-$ 1. Therefore, the first term of the left-hand side above is strictly greater than zero.
@@
$\sum_{j=a}^{n}\bigg{(}\mu\cdot\bigg{(}\sum_{i=1}^{j}\big{(}\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big{)}\\ \cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\\ +Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg{)}\bigg{)}\geq 0$ (9)

### v7

@@ Appendix D Proof that DReST-optimal policies are maximally USEFUL and maximally NEUTRAL
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)+\sum_{i=1}^{s-1}\big(\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1-i)\\ -\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-i)\big)\cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (3)
Here is how to interpret this equation. Selecting trajectory-length $x$ in mini-episode $e_{s}$ is guaranteed to yield at least $\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)$: the expected return that would be had if $x$ were selected in all s1s-1 previous mini-episodes. In addition, there is a probability of $Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-2\}$ that selecting $x$ in $e_{s}$ yields $\big(\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-2)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-1)\big)$: the extra expected return that would be had if $x$ were selected in only s2s-2 previous mini-episodes. In addition, there is a probability of $Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-3\}$ that selecting $x$ in $e_{s}$ yields $\big(\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-3)-\mathbb{E}_{\pi,e_{s}}(R|L=x,N_{e_{s}}(L=x)=s-2)\big)$: the extra expected return that would be had if $x$ were selected in only s3s-3 previous mini-episodes. And so on.
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x)=\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big(\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big)\cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}$ (4)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=y)=\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big(\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big)\cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}$ (5)
@@
$\mathbb{E}_{\pi,e_{s}}(R|L=x\lor L=y)=\\ Pr_{\pi,e_{s}}\{L=x\}\cdot\bigg(\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big(\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big)\cdot Pr_{\pi}\{N_{e_{s}}(L=x)\leq s-1-i\}\bigg)\\ +Pr_{\pi,e_{s}}\{L=y\}\cdot\bigg(\lambda^{s-1-\frac{s-1}{k}}+\sum_{i=1}^{s-1}\big(\lambda^{s-1-i-\frac{s-1}{k}}-\lambda^{s-i-\frac{s-1}{k}}\big)\cdot Pr_{\pi}\{N_{e_{s}}(L=y)\leq s-1-i\}\bigg)$ (6)
@@
(Note that $\pi_{n-1}$ implies one probability distribution over trajectory-lengths in the first n1n-1 mini-episodes $e_{1}$ to $e_{n-1}$ and implies a different probability distribution over trajectory-lengths in the final mini-episode $e_{n}$. Given that the environments in mini-episodes $e_{1}$ to $e_{n}$ are observationally-equivalent, policies like $\pi_{n-1}$ cannot be implemented. Nevertheless, it is useful to refer to policies like $\pi_{n-1}$ in proving Lemma D.2.)
Let $\pi_{n-2}$ be identical to $\pi_{n}$ except that $\pi_{n-2}$ selects trajectory-lengths $x$ and $y$ with the same probability $\mu$ in the final two mini-episodes $e_{n-1}$ to $e_{n}$. More precisely, $\pi_{n-2}$ is such that $Pr_{\pi_{n-2},e_{n-1}-e_{n}}\{L=x\}=Pr_{\pi_{n-2},e_{n-1}-e_{n}}\{L=y\}=\mu$. And so on.
@@
$\mathbb{E}_{\pi_{a-1},E}(R|L=x\lor L=y)=\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg)\\ +\mu\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg)\\ +\sum_{j=a}^{n}\bigg(\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg)\\ +\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg)\bigg)\\$ (7)
@@
$\mathbb{E}_{\pi_{a},E}(R|L=x\lor L=y)=\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg)\\ +(\mu-\Delta)\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg)\\ +\sum_{j=a}^{n}\bigg(\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg)\\ +\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg)\bigg)\\$ (8)
@@
$\mathbb{E}_{\pi_{a-1},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +\mu\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg)\\ +\mu\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a-1}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg)\\ +\sum_{j=a}^{n}\bigg(\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg)\\ +\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg)\bigg)\\ >\mathbb{E}_{\pi_{a},e_{1}-e_{a-1}}(R|L=x\lor L=y)\\ +(\mu+\Delta)\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\bigg)\\ +(\mu-\Delta)\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}\bigg)\\ +\sum_{j=a}^{n}\bigg(\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}\bigg)\\ +\mu\cdot\bigg(\lambda^{j-\frac{j}{k}}+\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\}\bigg)\bigg)\\$
@@
$\Delta\cdot\bigg(\lambda^{a-1-\frac{a-1}{k}}+\sum_{i=1}^{a-1}\big(\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}\big)\cdot(Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}-Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\})\bigg)\\ +\sum_{j=a}^{n}\bigg(\mu\cdot\bigg(\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}+Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}\\ -Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}-Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg)\bigg)>0$ (9)
By stipulation, $\Delta>0$. And since $0<\lambda<1$, $\lambda^{a-1-\frac{a-1}{k}}>0$ and $\lambda^{a-1-i-\frac{a-1}{k}}-\lambda^{a-i-\frac{a-1}{k}}>0$ for all $a$, $n$, and $k$. And since $Pr_{\pi_{a},e_{1}-e_{a}}\{L=x\}>Pr_{\pi_{a},e_{1}-e_{a}}\{L=y\}$, $Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}--Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}\geq 0$ for all $a$ and $i$ and $Pr_{\pi_{a}}\{N_{e_{a}}(L=y)\leq a-1-i\}--Pr_{\pi_{a}}\{N_{e_{a}}(L=x)\leq a-1-i\}>0$ for all $a$ and some $i$ such that $1\leq i\leq a-1$. Therefore, the first term of the left-hand side above is strictly greater than zero.
@@
$\sum_{j=a}^{n}\bigg(\mu\cdot\bigg(\sum_{i=1}^{j}\big(\lambda^{j-i-\frac{j}{k}}-\lambda^{j+1-i-\frac{j}{k}}\big)\cdot(Pr_{\pi_{a-1}}\{N_{e_{j}}(L=x)\leq j-i\}\\ +Pr_{\pi_{a-1}}\{N_{e_{j}}(L=y)\leq j-i\}-Pr_{\pi_{a}}\{N_{e_{j}}(L=x)\leq j-i\}-Pr_{\pi_{a}}\{N_{e_{j}}(L=y)\leq j-i\})\bigg)\bigg)\geq 0$ (10)